5 consecutive integers whose sum equals 70

Answer by pwac(253)

x= first integer
x+(x+1)+(x+2)+(x+3)+(x+4)=70 add left side
5x+10=70 subtract 10 from both sides
5x=60 divide both sides by 5
x=12
12+13+14+15+16=70
Pete